# Neutrino Oscillation Prospects with a Dual-Baseline Beam from BNL to SNOLAB and SURF

### v1

@@ 4 Probability
Table 2: The values of the oscillation parameters (taken from the global fit analysis in in 10.5281/zenodo.4726908 ; deSalas:2020pgw ) and their uncertainties used in our study. If the $3\sigma$ upper and lower limit of a parameter is $x_{u}$ and $x_{l}$ respectively, the $1\sigma$ uncertainty is $(x_{u}-x_{l})/3(x_{u}+x_{l})\%$ DUNE:2020ypp . The neutrino mass hierarchy was assumed to be normal.
In Fig. 2 (top, right panel) we have already shown $\text{P}_{\mu e}$ at 900 and 2900 km (calculated numerically using GLoBES by considering the best fit values of the oscillation parameters listed in Table 2). In order to understand the features of oscillation at different baselines, we analyze the probability expression in constant density for the $\nu_{\mu}\rightarrow\nu_{e}$ channel Akhmedov:2004ny .
@@
where $\alpha=\frac{\Delta m_{21}^{2}}{\Delta m_{31}^{2}}$ and $\Delta_{31}=\frac{\Delta m_{31}^{2}L}{4E}$. $A=\frac{2\sqrt{2}G_{F}N_{e}E}{\Delta m_{31}^{2}}$ is the matter effect where $N_{e}$ is the electron density along the neutrino propagation length. Assuming the earth matter to be electrically neutral (such that the ratio of the number density of electrons to that of nucleons is roughly 0.5), we can calculate the following quantity as Akhmedov:2004ny ,
@@
$L$ [km] $\rho\text{ [gm/cm}^{3}]$ $E^{\text{max}}_{1}$ [GeV] $E^{\text{max}}_{2}$ [GeV] $A\text{ at }E^{\text{max}}_{1}(E^{\text{max}}_{2})$
900 2.8 1.6 (1.8) 0.6 (0.6) 0.13 (0.05)
1300 2.95 2.1 (2.6) 0.8 (0.9) 0.19 (0.07)
2900 3.3 3.8 (5.9) 1.7 (2) 0.37 (0.17)
Table 3: This shows the line-averaged densities estimated using PREM profile Dziewonski:1981xy ; the energies $E^{\text{max}}_{1}$ and $E^{\text{max}}_{2}$ corresponding to the first and second oscillation maxima in matter (as well as in vacuum , - shown in parentheses); and the values of the matter effect term AA (see Eq. 3) corresponding to $E^{\text{max}}_{1}$ ($E^{\text{max}}_{2}$) for the three neutrino baselines.
@@
$n=1,2$ denote the first and second oscillation maxima (OM) respectively. The second term ($\propto\rho$) in the denominator of Eq. 5 comes from matter effect. Using the average densities ($\rho$) of the three baselines (following the Preliminary Reference Earth Model (PREM) Dziewonski:1981xy ), the values of $E^{\text{max}}_{1}$ and $E^{\text{max}}_{2}$ are estimated in Table 3. It is evident that, for longer baselines, the oscillation maxima shift to higher energies, thereby making them relatively easier to probe. The estimated values of $E^{\text{max}}_{1}$ and $E^{\text{max}}_{2}$ for 2900 (900) km are approximately 3.8 (1.6) GeV and 1.7 (0.6) GeV respectively. This suggests that the EIC-SURF 2900 km baseline could be a very promising option, as it allows the possibility of observing more than one oscillation cycle. Furthermore, the presence of the overall factor (1A)(1-A) in the denominator of Eq. 4 increases the amplitude of oscillation with energy and matter density, - which in turn increases with baseline (see Table 3 for an estimate of the factor $A$ at first and second oscillation maxima). Thus the magnitude of $\text{P}_{\mu e}$ at 2900 km is almost twice than at 900 km around the first oscillation maximum (as evident from Fig. 2: top row, 3rd column).
@@
Figure 3: The solid curves show the quantities $\Delta P_{\mu e}^{\text{I}},\Delta P_{\mu e}^{\text{II}}$ and $\Delta P_{\mu e}=(\Delta P_{\mu e}^{\text{I}}+\Delta P_{\mu e}^{\text{II}})$ from Eq. 4 as functions of energy for the three baselines (900 km, 1300 km, 2900 km). The dotted red curves indicate the intrinsic CP asymmetry $(\Delta P_{\mu e}-\Delta P_{\mu e}^{\delta_{\text{CP}}=0})$ originating only from the phase $\delta_{\text{CP}}$ and free from matter effects. The pair of black dotted vertical lines in each panel indicates the relevant $E^{\text{max}}_{1}$ and $E^{\text{max}}_{2}$.
@@
In this context, it is worth noting that for large baselines (e.g., 2900 km), matter effects become significant and must be carefully disentangled from genuine CP violation effects Arafune:1997hd ; Nunokawa:2007qh . In such cases, a more appropriate way to quantify the intrinsic CP asymmetry arising from the phase $\delta_{\text{CP}}$ is to subtract the asymmetry induced purely by matter effects in the absence of the CP violating phase $\delta_{\text{CP}}$. This is expressed as $(\Delta P_{\mu e}-\Delta P_{\mu e}^{\delta_{\text{CP}}=0})$ Ohlsson:2013ip ; Rout:2017udo .
@@
In order to analyze the CP asymmetries when $\delta_{\text{CP}}$ is varied, we show the heatplots of $\Delta P_{\mu e}$ and $(\Delta P_{\mu e}-\Delta P_{\mu e}^{\delta_{\text{CP}}=0})$ (both calculated numerically using GLoBES) in the plane of $E-\delta_{\text{CP}}$ in Fig. 4 for the three baselines. The red (blue) color indicates a positive (negative) value while the saturation of the color signifies the magnitude of the asymmetry. As expected, the magnitude of the asymmetry is highest around the maximally CP violating values $\delta_{\text{CP}}\simeq\pm\pi/2$. For simplicity, the behaviour of $\Delta P_{\mu e}$ for small matter effects can be understood from the vacuum approximation (Eq. 8). The $\sin\delta_{\text{CP}}$ factor with a minus sign in Eq. 8 explains why $\Delta P_{\mu e}$ is negative (positive) for $\delta_{\text{CP}}>0$ ($\delta_{\text{CP}}<0$). Such antisymmetry around $\delta_{\text{CP}}\simeq 0$ gets distorted by matter effects,- most prominently for 2900 km as illustrated in Fig. 4. The bottom row of Fig. 4 illustrates the intrinsic asymmetries largely free from matter effects and with less distorted patterns.
The heatmaps clearly show that as the baseline increases, multiple oscillation cycles can, in principle, be probed. This, in turn, leads to enhanced CP asymmetry across the accessible energy range. To get a simple numerical estimate of the number of oscillation cycles present within a given energy interval for a fixed $\delta_{\text{CP}}$, we use the vacuum approximation (Eq. 8) for brevity. The number of cycles is determined by the argument of the term $\sin^{2}\Delta_{31}=(1-\cos 2\Delta_{31})/2$. Since $2\Delta_{31}\simeq 10^{-3}\times L[\text{km}]/E[\text{GeV}]$, the number of cycles $N_{\text{cycles}}$ in the energy interval $E_{1},E_{2}]$ (in GeVs) is given by,
@@
Thus in the range [0.210][0.2-10] GeV, the estimated number of oscillation cycles for the three baselines (900, 1300, 2900 km) are 4.4, 6.4, 14.2 respectively. This approximate analytical estimate is in good agreement with the numerical simulation, as reflected by the number of red or blue blobs observed in each panel of Fig. 4.
Figure 5: Electron neutrino appearance events at THEIA for the EIC to SURF baseline (2900 km). Here we show only signals (dotted) as well as events (signal + background).
@@ 5 Event spectra
Figure 6: Muon neutrino disappearance events at THEIA for the EIC to SURF baseline (2900 km)
@@
We assume a 7 yrs. of runtime (3.5 yrs. each in neutrino and antineutrino mode) and a water based liquid scintillator or WbLS  [^3] WbLS can detect particles based on both scintillation and Cherenkov radiation, the latter providing a directional capability as well OrebiGann:2015gus ; Fischer:2018zsr ; Kaptanoglu:2019gtg ; Callaghan:2023oyu . WbLS is a proposed detector in the THEIA collaboration Theia:2019non ; Theia:2022uyh ; mary_dune_theia_2023 . WbLS detectors can help in CP violation studies Ge:2022iac ; in antineutrino detection Zsoldos:2022mre ; in low energy astrophysical neutrino detection DeGouvea:2020ang ; Sawatzki:2020mpb ; DeRomeri:2021xgy ; Borboruah:2025hai ; in neutrino-nucleon interaction studies ANNIE:2023yny ; as well as in studying rare processes such as neutrino decay or absolute neutrino mass DeGouvea:2020ang ; Parker:2023cos ; Chauhan:2022wgj . with a fiducial mass of 17 kt for both the baselines. Following the discussions in Theia:2019non , regarding the capability of WbLS to identify Cherenkov rings, we implement separate channels in the corresponding .glb file of GLoBES for $\nu_{e}$} signals: one, two or three-ring rules with either zero or one Michel electrons from stopped pion or muon decay. The channels with $n$ Cherenkov rings and $i$ Michel electrons are referred to as $nRiD$ with $n=1,2,3$ and $i=0,1$ in this case. We further implement different channels based on whether the $\nu_{e}$} signal is charge-current quasi-elastic (CCQE) or charge-current non-quasielastic (CCNQE).
@@
For the $\nu_{\mu}$} signals we implement two separate channels, namely $1R0D,1R1D$ and consider charge current interactions inclusively (CC), i.e., both CCQE and CCNQE together in one channel in GLoBES. The single ring channels provide the dominant contributions to the $\nu_{e}$} and $\nu_{\mu}$} event spectra. All these channels include pre and post smearing efficiencies taken from guang_theia_2018 ; Theia:2019non . For instance, in the energy range of 1-5 GeV, the pre (post) smearing efficiencies for the channel $1R0D$ for CCQE $\nu_{e}$} signal lie around $75\%-85\%$ ($85\%-95\%$). For $\nu_{\mu}$} channels, we have used an overall flat efficiency of $90\%$ throughout the entire energy range. For the same $1R0D$ channel, the migration matrices for smearing in the WbLS detector correspond to an energy resolution ($\sigma/E$) of $6\%-15\%$ for CCQE $\nu_{e}$} signal in 1-5 GeV energy range. For the $\nu_{\mu}$} channel, we have assumed a Gaussian smearing of $10\%$. In order to understand the impact of smearing, we have also estimated the event spectra by assuming a detector with perfect energy resolution ($\it i.e.$}, no energy smearing).
@@
Figure 7: Electron neutrino appearance events at WbLS detector for the EIC to SNOlab baseline (900 km)
Figure 8: Muon neutrino disappearance events at WbLS detector for the EIC to SNOlab baseline (900 km)
@@ 6 $\Delta\chi^{2}$ Analysis
where the index $y$ is summed over the energy bins in the range 0.5180.5-18 GeV  [^4] We have a total of $65$ energy bins in the range 0.5180.5-18 GeV: $60$ bins each having a width of $0.125$ GeV in the energy range of 0.580.5-8 GeV and $5$ bins with 2 GeV width in 8-18 GeV. Also note that, the analysis range for the SURF (SNO) site was taken as 0.8180.8-18 (2.218)(2.2-18) GeV for the case where only the first oscillation maximum was consider by ignoring the second. unless otherwise mentioned explicitly. The index $j$ corresponds to the six appearance channels nRiDnRiD ($n$ Cerenkov rings and $i$ Michel electrons from stopped pion or muon decay) with $n=1,2,3$, and $i=0,1$, as discussed in Sec. 5. Finally the index $x$ runs over the neutrino and antineutrino modes. $N^{\text{data}}$ corresponds to the set of events generated for any value of the CP phase $\delta_{\text{CP}}\in[-\pi,\pi]$ with the other oscillation parameters fixed to the global best fit values. $N^{\text{fit}}$ corresponds to the events generated only for the CP conserving values of the phase: $\delta_{\text{CP}}^{\text{fit}}=0,\pi$ with other oscillation parameters being varied in their allowed range (see Table 2 for both the bestfit values and the range of variation of the oscillation parameters). The terms in the first row of the right-hand-side of Eq. (6) correspond to the statistical contribution. The first two terms in the statistical contribution to the $\Delta\chi^{2}$ constitute the algebraic difference ($N^{\text{fit}}-N^{\text{data}}$) while the last term (i.e., the logarithmic-term) corresponds to the fractional difference between the two sets of events [^5] Note that the definition of $\Delta\chi^{2}$ described in Eq. 6 is Poissonian in nature. In the limit of large number of events (N), this reduces to the Gaussian form : $\lim_{N\to\infty}\Delta\chi^{2}(\delta_{\text{CP}})\simeq{\text{min}}\Bigg[\sum_{x}^{\text{mode}}\sum_{j}^{\text{channel}}\sum_{i}^{\text{bin}}\frac{\Big\{N_{ijx}^{\text{fit}}(\delta_{\text{CP}}=0,\pi)-N_{ijx}^{\text{data}}(\delta_{\text{CP}})\Big\}^{2}}{N_{ijx}^{\text{data}}(\delta_{\text{CP}})}+\text{prior}+\text{systematics}\Bigg].$ .
@@
Figure 9: This shows the CP violation sensitivities achieved from the neutrino beam (produced using 1 MW proton beam at BNL) at a 17 kt WbLS detector at SURF (left panel), and at SNO (right panel) site for a $(3.5+3.5)$ yrs. of $(\nu+\bar{\nu})$ mode runtimes. The red (green) curves correspond to a detector with a realistic energy smearing (without smearing). The sensitivities shown by the dotted curves were estimated using a curtailed energy range so as to consider only the first oscillation maximum, while the solid and dashed curves additionally consider second and higher oscillation maxima as well.
The degree of this deviation is controlled by $\sigma_{pl}$ which is the uncertainty in the prior measurement of the best-fit values of $p^{\text{data}}$ (see the last column of Table 2 for the values used in the present analysis). The systematics-term accounts for the variation of the systematic/nuisance parameters. $\eta$ is the set of values of $k$-systematics parameters $\{\eta_{1}$, $\eta_{2}$, $\dots$ $\eta_{k}\}$ (which also impact $N^{\text{fit}}$ as normalization uncertainties) while $\sigma_{k}$ is the uncertainty in the corresponding systematics. This way of treating the nuisance parameters in the $\Delta\chi^{2}$ calculation is known as the method of pulls Huber:2002mx ; Fogli:2002pt ; GonzalezGarcia:2004wg ; Gandhi:2007td . Regarding the systematics Theia:2019non , the $\nu_{e}$ and $\bar{\nu}_{e}$ signal modes have a normalization uncertainty of $2\%$ each. As discussed in Sec. 5, backgrounds were neglected in the $\Delta\chi^{2}$ analysis.
The final estimate of the $\Delta\chi^{2}$ as a function of the CP phase $\delta_{\text{CP}}$ is obtained after minimizing the entire quantity within the square bracket in Eq. 6 over the CP conserving values of the phase ($\delta_{\text{CP}}^{\text{fit}}=0,\pi$); the fit oscillation parameters in their allowed range; as well as over the systematics $\eta$. This minimization is also referred to as marginalization. Technically, this procedure is the frequentist method of hypotheses testing Fogli:2002pt ; Qian:2012zn .

### v2

@@ 4 Probability
Table 2: The values of the oscillation parameters (taken from the global fit analysis in P. F. De Salas et al. (2021); de Salas et al. (2021)) and their uncertainties used in our study. If the $3\sigma$ upper and lower limit of a parameter is $x_{u}$ and $x_{l}$ respectively, the $1\sigma$ uncertainty is $(x_{u}-x_{l})/3(x_{u}+x_{l})\%$ Abi and others (2020b). The neutrino mass hierarchy was assumed to be normal.
In Fig. 2 (top, right panel) we have already shown $\text{P}_{\mu e}$ at 900 and 2900 km (calculated numerically using GLoBES by considering the best fit values of the oscillation parameters listed in Table 2). In order to understand the features of oscillation at different baselines, we analyse the probability expression in constant density for the $\nu_{\mu}\rightarrow\nu_{e}$ channel Akhmedov et al. (2004).
@@
where $\alpha=\frac{\Delta m_{21}^{2}}{\Delta m_{31}^{2}}$ and $\Delta_{31}=\frac{\Delta m_{31}^{2}L}{4E}$. $A=\frac{2\sqrt{2}G_{F}N_{e}E}{\Delta m_{31}^{2}}$ is the matter effect where $N_{e}$ is the electron density along the neutrino propagation length. For Earth matter, which is mainly composed of medium to heavy nuclei (Si, Fe, Mg etc.), the average value of the ratio of the number density of electrons to that of nucleons is roughly 0.5. We can then calculate the following quantity as Akhmedov et al. (2004),
@@
$L$ $\rho$ $E^{\text{max}}_{1}$ [GeV] $E^{\text{max}}_{2}$ [GeV] $A\text{ at }E^{\text{max}}_{1}(E^{\text{max}}_{2})$
[km] $\text{ [gm/cm}^{3}]$ Approx. Exact Approx. Exact Approx. Exact
(Eq. 5) (numerically)
900 2.8 1.6 (1.8) 1.6 (1.8) 0.6 (0.6) 0.6 (0.6) 0.13 (0.05) 0.13 (0.05)
1300 2.95 2.1 (2.6) 2.2 (2.6) 0.8 (0.9) 0.8 (0.9) 0.19 (0.07) 0.20 (0.07)
2900 3.3 3.8 (5.9) 4.1 (6.1) 1.7 (2) 1.8 (2) 0.37 (0.17) 0.40 (0.18)
Table 3: This shows the line-averaged densities ($\rho$) estimated using PREM profile Dziewonski and Anderson (1981) for the three baselines ($L$); the energies $E^{\text{max}}_{1}$ and $E^{\text{max}}_{2}$ corresponding to the first and second oscillation maxima in matter (as well as in vacuum , - shown in parentheses); and the values of the matter effect term AA (see Eq. 3) corresponding to $E^{\text{max}}_{1}$ ($E^{\text{max}}_{2}$) for the three neutrino baselines. For $E^{\text{max}}_{1,2}$ and $A$, both the approximate values (from Eq. 5) and the exact numerical values are tabulated for comparison.
@@
$n=1,2$ denote the first and second oscillation maxima (OM) respectively. The second term ($\propto\rho$) in the denominator of Eq. 5 comes from matter effect. Using the average densities ($\rho$) of the three baselines (following the Preliminary Reference Earth Model (PREM) Dziewonski and Anderson (1981)  [^4] Note that PREM profile does not account for local crustal variations due to topography or regional geological structures. However, these local effects are not expected to have any significant impact on the results presented here, and the use of the PREM is therefore adequate for our purposes.), the values of $E^{\text{max}}_{1}$ and $E^{\text{max}}_{2}$ are estimated in Table 3. It is evident that, for longer baselines, the oscillation maxima shift to higher energies, thereby making them relatively easier to probe. The estimated values of $E^{\text{max}}_{1}$ and $E^{\text{max}}_{2}$ for 2900 (900) km are approximately 3.8 (1.6) GeV and 1.7 (0.6) GeV respectively. This suggests that the EIC-SURF 2900 km baseline could be a very promising option, as it allows the possibility of observing more than one oscillation cycle. Furthermore, the presence of the overall factor (1A)(1-A) in the denominator of Eq. 4 increases the amplitude of oscillation with energy and matter density, - which in turn increases with baseline (see Table 3 for an estimate of the factor $A$ at first and second oscillation maxima). Thus the magnitude of $\text{P}_{\mu e}$ at 2900 km is almost twice than at 900 km around the first oscillation maximum (as evident from Fig. 2: top row, 3rd column).
@@
Figure 3: The solid curves show the quantities $\Delta P_{\mu e}^{\text{I}},\Delta P_{\mu e}^{\text{II}}$ and $\Delta P_{\mu e}=(\Delta P_{\mu e}^{\text{I}}+\Delta P_{\mu e}^{\text{II}})$ from Eq. 4 as functions of energy for the three baselines (900 km, 1300 km, 2900 km). The dotted red curves indicate the intrinsic CP asymmetry $(\Delta P_{\mu e}-\Delta P_{\mu e}^{\delta_{\text{CP}}=0})$ originating only from the phase $\delta_{\text{CP}}$ and free from matter effects. For comparison, we have also included the exact values of $\Delta P_{\mu e}$ calculated numerically using GLoBES. The pair of black dotted vertical lines in each panel indicates the relevant $E^{\text{max}}_{1}$ and $E^{\text{max}}_{2}$.
@@
In this context, it is worth noting that for large baselines (e.g., 2900 km), matter effects become significant and must be carefully disentangled from genuine CP violation effects Arafune et al. (1997); Nunokawa et al. (2008). In such cases, a more appropriate way to quantify the intrinsic CP asymmetry arising from the phase $\delta_{\text{CP}}$ is to subtract the asymmetry induced purely by matter effects in the absence of the CP violating phase $\delta_{\text{CP}}$. This is expressed as $(\Delta P_{\mu e}-\Delta P_{\mu e}^{\delta_{\text{CP}}=0})$ Ohlsson et al. (2013); Rout et al. (2017).
@@
In order to analyse the CP asymmetries when $\delta_{\text{CP}}$ is varied, we show the heatplots of $\Delta P_{\mu e}$ and $(\Delta P_{\mu e}-\Delta P_{\mu e}^{\delta_{\text{CP}}=0})$ (both calculated numerically using GLoBES) in the plane of $E-\delta_{\text{CP}}$ in Fig. 4 for the three baselines. The red (blue) colour indicates a positive (negative) value while the saturation of the colour signifies the magnitude of the asymmetry. As expected, the magnitude of the asymmetry is highest around the maximally CP violating values $\delta_{\text{CP}}\simeq\pm\pi/2$. For simplicity, the behaviour of $\Delta P_{\mu e}$ for small matter effects can be understood from the vacuum approximation (Eq. 8). The $\sin\delta_{\text{CP}}$ factor with a minus sign in Eq. 8 explains why $\Delta P_{\mu e}$ is negative (positive) for $\delta_{\text{CP}}>0$ ($\delta_{\text{CP}}<0$). Such antisymmetry around $\delta_{\text{CP}}\simeq 0$ gets distorted by matter effects,- most prominently for 2900 km as illustrated in Fig. 4. The bottom row of Fig. 4 illustrates the intrinsic asymmetries largely free from matter effects and with less distorted patterns.
The heatmaps clearly show that as the baseline increases, multiple oscillation cycles can, in principle, be probed. This, in turn, leads to enhanced CP asymmetry across the accessible energy range. To get a simple numerical estimate of the number of oscillation cycles present within a given energy interval for a fixed $\delta_{\text{CP}}$, we use the vacuum approximation (Eq. 8) for brevity. The number of cycles is determined by the argument of the term $\sin^{2}\Delta_{31}=(1-\cos 2\Delta_{31})/2$. Since $2\Delta_{31}\simeq 2\pi\times 10^{-3}\times L[\text{km}]/E[\text{GeV}]$, the number of cycles $N_{\text{cycles}}$ in the energy interval $E_{1},E_{2}]$ (in GeVs) is given by,
@@
Thus in the range [0.210][0.2-10] GeV, the estimated number of oscillation cycles for the three baselines (900, 1300, 2900 km) are 4.4, 6.4, 14.2 respectively. This approximate analytical estimate is in good agreement with the numerical simulation, as reflected by the number of red or blue blobs observed in each panel of Fig. 4. In this context, it needs to be mentioned that counting the number of oscillation cycles in actual data using Eq. 9 requires both very good energy resolution (especially at lower energies), as well as small uncertainties in the neutrino flux (i.e., less systematics). Nevertheless, Eq. 9 can potentially give an instant, semi-quantitative insight into the oscillatory nature of $\nu_{\mu}\rightarrow\nu_{e}$ transition for a given baseline and energy interval.
Figure 5: $\nu_{e}$} appearance events (signals only) at a 17 kt WbLS detector for the EIC to SURF baseline (2900 km). The blue curve indicates events for a WbLS detector with a realistic energy smearing, while the red curves show the case of no energy smearing. The light yellow regions show the underlying $\text{P}_{\mu e}$ (scaled with a constant arbitrary factor for better visibility).
@@ 5 Event spectra
Figure 6: $\nu_{\mu}$} disappearance events (signals only) at a 17 kt WbLS detector for the EIC to SURF baseline (2900 km). The blue curve indicates events for a WbLS detector with a realistic energy smearing, while the red curves show the case of no energy smearing. The light yellow regions show the underlying $\text{P}_{\mu e}$ (scaled).
@@
We assume a 7 yrs. of runtime (3.5 yrs. each in neutrino and antineutrino mode) and a water based liquid scintillator or WbLS  [^5] WbLS can detect particles based on both scintillation and Cherenkov radiation, the latter providing a directional capability as well Orebi Gann (2015); Fischer (2018); Kaptanoglu et al. (2020); Callaghan et al. (2023). WbLS is a proposed detector in the THEIA collaboration Askins and others (2020, 2022); Mary Bishai (DUNE collaboration) (2023). WbLS detectors can help in CP violation studies Ge et al. (2022); in antineutrino detection Zsoldos et al. (2022); in low energy astrophysical neutrino detection De Gouvêa et al. (2020); Sawatzki et al. (2021); De Romeri et al. (2021); Borboruah et al. (2025); in neutrino-nucleon interaction studies Ascencio-Sosa and others (2024); as well as in studying rare processes such as neutrino decay or absolute neutrino mass De Gouvêa et al. (2020); Parker and Wurm (2024); Chauhan (2022). with a fiducial mass of 17 kt for both the baselines. Following the discussions in Askins and others (2020), regarding the capability of WbLS to identify Cherenkov rings, we implement separate channels in the corresponding .glb file of GLoBES for $\nu_{e}$} signals: one, two or three-ring rules with either zero or one Michel electrons from stopped pion or muon decay. The channels with $n$ Cherenkov rings and $i$ Michel electrons are referred to as $nRiD$ with $n=1,2,3$ and $i=0,1$ in this case. We further implement different channels based on whether the $\nu_{e}$} signal is charge-current quasi-elastic (CCQE) or charge-current non-quasielastic (CCNQE).
@@
For the $\nu_{\mu}$} signals we implement two separate channels, namely $1R0D,1R1D$ and consider charge current interactions inclusively (CC), i.e., both CCQE and CCNQE together in one channel in GLoBES. The single ring channels provide the dominant contributions to the $\nu_{e}$} and $\nu_{\mu}$} event spectra. All these channels include pre and post smearing efficiencies taken from Guang Yang (2018); Askins and others (2020). For instance, in the energy range of 1-5 GeV, the pre (post) smearing efficiencies for the channel $1R0D$ for CCQE $\nu_{e}$} signal lie around $75\%-85\%$ ($85\%-95\%$). For $\nu_{\mu}$} channels, we have used an overall flat efficiency of $90\%$ throughout the entire energy range. For the same $1R0D$ channel, the migration matrices for smearing in the WbLS detector correspond to an energy resolution ($\sigma/E$) of $6\%-15\%$ for CCQE $\nu_{e}$} signal in 1-5 GeV energy range. For the $\nu_{\mu}$} channel, we have assumed a Gaussian smearing of $10\%$. In order to understand the impact of smearing, we have also estimated the event spectra by assuming a detector with perfect energy resolution ($\it i.e.$}, no energy smearing).
@@
Figure 7: $\nu_{e}$} appearance events (signals only) at WbLS detector for the EIC to SNOlab baseline (900 km).
@@ 6 $\Delta\chi^{2}$ Analysis
where the index $y$ is summed over the energy bins in the range 0.5180.5-18 GeV  [^6] We have a total of $65$ energy bins in the range 0.5180.5-18 GeV: $60$ bins each having a width of $0.125$ GeV in the energy range of 0.580.5-8 GeV and $5$ bins with 2 GeV width in 8-18 GeV. Also note that, the analysis range for the SURF (SNO) site was taken as 0.8180.8-18 (2.218)(2.2-18) GeV for the case where only the first oscillation maximum was consider by ignoring the second. unless otherwise mentioned explicitly. The index $j$ corresponds to the six appearance channels nRiDnRiD ($n$ Cerenkov rings and $i$ Michel electrons from stopped pion or muon decay) with $n=1,2,3$, and $i=0,1$, as discussed in Sec. 5. Finally the index $x$ runs over the neutrino and antineutrino modes. $N^{\text{data}}$ corresponds to the set of events generated for any value of the CP phase $\delta_{\text{CP}}\in[-\pi,\pi]$ with the other oscillation parameters fixed to the global best fit values. $N^{\text{fit}}$ corresponds to the events generated only for the CP conserving values of the phase: $\delta_{\text{CP}}^{\text{fit}}=0,\pi$ with other oscillation parameters being varied in their allowed range (see Table 2 for both the best fit values and the range of variation of the oscillation parameters). The terms in the first row of the right-hand-side of Eq. (6) correspond to the statistical contribution.
Figure 8: $\nu_{\mu}$} disappearance events (signals only) at WbLS detector for the EIC to SNOlab baseline (900 km).
The first two terms in the statistical contribution to the $\Delta\chi^{2}$ constitute the algebraic difference ($N^{\text{fit}}-N^{\text{data}}$) while the logarithmic-term proportional to $\ln(N^{\text{data}}/N^{\text{fit}})$ takes into account the relative difference or ratio of the two sets of events [^7] Note that the definition of $\Delta\chi^{2}$ described in Eq. 6 is Poissonian in nature. In the limit of large number of events (N), this reduces to the Gaussian form : $\lim_{N\to\infty}\Delta\chi^{2}(\delta_{\text{CP}})\simeq{\text{min}}\Bigg[\sum_{x}^{\text{mode}}\sum_{j}^{\text{channel}}\sum_{i}^{\text{bin}}\frac{\Big\{N_{ijx}^{\text{fit}}(\delta_{\text{CP}}=0,\pi)-N_{ijx}^{\text{data}}(\delta_{\text{CP}})\Big\}^{2}}{N_{ijx}^{\text{data}}(\delta_{\text{CP}})}+\text{prior}+\text{systematics}\Bigg].$ .
@@
Figure 9: This shows the CP violation sensitivities achieved from the neutrino beam (produced using 1 MW proton beam at BNL) at a 17 kt WbLS detector at SURF (left panel), and at SNO (right panel) site for a $(3.5+3.5)$ yrs. of $(\nu+\bar{\nu})$ mode runtimes. The sensitivities shown by the dotted curves were estimated using a curtailed energy range so as to consider only the first oscillation maximum, while the solid and dashed curves additionally consider second and higher oscillation maxima as well.
The degree of this deviation is controlled by $\sigma_{pl}$ which is the uncertainty in the prior measurement of the best-fit values of $p^{\text{data}}$ (see the last column of Table 2 for the values used in the present analysis). The systematics-term accounts for the variation of the systematic/nuisance parameters. $\eta$ is the set of values of $k$-systematics parameters $\{\eta_{1}$, $\eta_{2}$, $\dots$ $\eta_{k}\}$ (which also impact $N^{\text{fit}}$ as normalization uncertainties) while $\sigma_{k}$ is the uncertainty in the corresponding systematics. This way of treating the nuisance parameters in the $\Delta\chi^{2}$ calculation is known as the method of pulls Huber et al. (2002); Fogli et al. (2002); Gonzalez-Garcia and Maltoni (2004); Gandhi et al. (2007). Regarding the systematics Askins and others (2020), the $\nu_{e}$ and $\bar{\nu}_{e}$ signal modes have a normalization uncertainty of $2\%$ each. As discussed in Sec. 5, backgrounds were neglected in the $\Delta\chi^{2}$ analysis.
The final estimate of the $\Delta\chi^{2}$ as a function of the CP phase $\delta_{\text{CP}}$ is obtained after minimizing the entire quantity within the square bracket in Eq. 6 over the CP conserving values of the phase ($\delta_{\text{CP}}^{\text{fit}}=0,\pi$); the fit oscillation parameters in their allowed range; as well as over the systematics $\eta$. This minimization is also referred to as marginalization. Technically, this procedure is the frequentist method of hypotheses testing Fogli et al. (2002); Qian et al. (2012).